# Partially Stochastic Infinitely Deep Bayesian Neural Networks

### v1

@@ Abstract
In this paper, we present Partially Stochastic Infinitely Deep Bayesian Neural Networks, a novel family of architectures that integrates partial stochasticity into the framework of infinitely deep neural networks. Our new class of architectures is designed to improve the limitations of existing architectures around computational efficiency at training and inference time. To do this, we leverage the advantages of partial stochasticity in the infinite-depth limit which include the benefits of full stochasticity e.g. robustness, uncertainty quantification, and memory efficiency, whilst improving their limitations around computational efficiency at training and inference time. We present a variety of architectural configurations, offering flexibility in network design including different methods for weight partition. We also provide mathematical guarantees on the expressivity of our models by establishing that our network family qualifies as Universal Conditional Distribution Approximators. Lastly, empirical evaluations across multiple tasks show that our proposed architectures achieve better downstream task performance and uncertainty quantification than their counterparts while being significantly more efficient.
@@ 1 Introduction
To overcome these challenges, previous research has explored partially stochastic networks as a potential solution in finite-depth settings (Daxberger et al. 2021b; Kristiadi et al. 2020). These networks restrict stochasticity to a fraction of the model’s parameters, i.e. $\Theta=\Theta_{S}\cup\Theta_{D}$, where one learns both a distribution over $\Theta_{S}$ and point estimates for $\Theta_{D}$. These methods aim to achieve greater computational efficiency while retaining the benefits of Bayesian approaches. Initially, the main idea of these architectures was to approximate fully stochastic networks with a trade-off between computational efficiency and accuracy. However, Sharma et al. (2023) provide both theoretical and empirical evidence to demonstrate that partially stochastic networks can achieve comparable, if not superior, performance to their fully stochastic counterparts. As there is no trivial way to split the weights into subsets of deterministic and stochastic weights in the infinite-depth limit, previous research has only explored finite-depth regimes. However, studies such as Farquhar et al. (2020) hint towards the viability of extending this approach to an infinite-depth configuration.
@@ 2 Background and Preliminaries
In Bayesian Neural Networks (Lampinen & Vehtari 2001; Titterington 2004), parameters are treated as random variables, with the aim to infer their posterior distribution $p(w|\mathcal{D})$ given a dataset $\mathcal{D}$ and prior $p(w)$. Approximate inference involves finding an approximate posterior, often through a parametrisation $q_{\theta}(w)$. Specifically, variational inference methods intend to maximise the Evidence Lower Bound (ELBO), or equivalently, minimise the Kullback-Leibler (KL) divergence between $q_{\theta}(w)$ and $p(w|\mathcal{D})$. The gradient of this objective can be computed using Monte-Carlo samples via Stochastic Variational Inference (Hoffman et al., 2013). In addition, previous studies suggest that full stochasticity is unnecessary for effective posterior approximation in finite-depth BNNs (Sharma et al., 2023). Extending these findings, we show that partial stochasticity preserves the expressivity of BNNs in the infinite-depth regime.
@@
where $\mu:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ (the drift), $\sigma:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times m}$ (the volatility or diffusion matrix) are Lipschitz functions in the state, and satisfy a linear growth condition. Such conditions ensure the existence and uniqueness of a strong solution, given a square-integrable initial condition $X_{0}$. In infinite-depth Bayesian neural networks, Xu et al. (2022) propose to model the prior on weights $(w_{t})_{t\in[0,1]}$ as a SDE
@@
and parametrise the approximate posterior using a Neural SDE, where the drift $f_{q}$ is a neural network with parameters $\theta$
@@
We use the following infinite-dimensional version of the KL divergence (Tzen & Raginsky 2019; Li et al. 2020)
@@
with $\mu_{p},\mu_{q}$ being probability measures on the path space $\mathcal{P}\subset\mathbb{R}^{[0,1]}$ induced by (1) and (2) respectively. We show that when only a fraction of the weights are stochastic, such approximate posteriors yield UCDAs.
@@ 3 Method
where $t$ is the layer index, $h_{t}\in\mathbb{R}^{d_{h}}$ represents a vector of hidden unit activations at layer $t$, the input is $h_{0}=x$, $w_{t}$ are the parameters for layer t, and $\delta=1$. Taking $\delta=1/T$ and letting $T\rightarrow\infty$ yields a Neural ODE $\mathrm{d}h_{t}=f_{h}(t,h_{t};w_{t})\mathrm{d}t$, where $f_{h}$ is a neural network with parameters $w_{t}$. These parameters can also be modelled by a neural network (more precisely, a hypernetwork), which in the infinite-depth limit becomes itself a neural differential equation $\mathrm{d}w_{t}=f_{q}(t,w_{t};\theta)\mathrm{d}t$ (or, in the Bayesian setting, a Neural SDE). Therefore, $f_{q}$ is a neural network parametrised by $\theta$ which models the continuous evolution of the parameters of $f_{h}$. Finally, in this framework, the only trainable parameters are the vector $\theta$ and $w_{0}$.
@@ 3.1 Vertical separation of the weights
Notice that the diffusion of the approximate posterior is fixed to match the prior diffusion. This is to ensure that the KL divergence between the prior and posterior measures on paths $\mu_{p}$ and $\mu_{q}$ respectively, is finite. As this is mentioned without proof by Li et al. (2020) and Xu et al (2022), we put it as a lemma and give a detailed proof in Appendix A.
@@
In conclusion. the dynamics of the whole system can then be written as a single SDE in $\mathbb{R}^{d_{h}+d_{w}}$
@@ 4.2 Partially Stochastic Infinitely Deep Bayesian Neural Networks are UCDAs
${}^{\dagger}$ 99.46 $\pm$ 0.00 2.88 ±$\pm 0.94$ 87.35 ±$\pm 0.00$ 8.47 $\pm$ 0.39
${}^{\dagger}$ 98.90 ±$\pm 0.04$ 1.11 $\pm$ 0.10 88.30 $\pm$ 0.29 8.71 ±$\pm 0.21$
MFVI ResNet ${}^{\dagger}$ 99.40 $\pm$ 0.00 2.76 $\pm$ 1.28 86.97 $\pm$ 0.00 3.04 $\pm$ 0.94
${}^{\dagger}$ 86.48 1.95
@@
Deep ${}^{\dagger}$ 89.22 2.79
HMC (“gold ${}^{\dagger}$ 98.31 1.79 90.70 5.94
MFVI ${}^{\dagger}$ 98.81 ±$\pm 0.00$ 2.63 ±$\pm 0.31$ 81.59 ±$\pm 0.01$ 3.62 ±$\pm 0.40$
MFVI ${}^{\dagger}$ 98.77 ±$\pm 0.01$ 2.82 ±$\pm 1.34$ 80.62 ±$\pm 0.00$ 4.29 ±$\pm 1.10$
@@ 5.1 Experimental Setup
[…]
@@ 5.3 Uncertainty Quantification
In the discrete neural networks domain, partial stochasticity is initially conceptualised as a simplified approximation of fully stochastic neural networks aimed at providing just enough uncertainty quantification. However, as we have theoratically established in earlier sections, partial stochasticity provides an expressive framework to approximate any posterior distribution. In this section, we further prove this empirically by looking at the expected calibration error (ECE) of the models at test time, and by performing predictive entropy analysis of out-of-distribution (OOD) and in-distribution (ID) samples.
[…]
@@ 5.4 Efficiency and Learning Rate
Table 2: Models’ average time per training epoch and at inference.
Model Epoch Time (s) Inference Time (s)
@@
ODEFirst 289.4   31.7
SDEFirst $295.1$   31.7
Next, we explore the learning rate of our models, i.e. how many epochs the network takes to get to acceptable downstream task performance. PSDE-BNN takes considerably less time to achieve the reported classification accuracies. While our models trained for 100 epochs, SDE-BNN (and as reported by Xu et al. 2022) required 300 epochs. The evolution after the 80% test accuracy for SDE-BNN was especially long. This indicates that PSDE-BNNs were both faster at each epoch, and only needed a third of the epochs to achieve similar (although still better) downstream performance. Therefore, the overall training time was 74.1% faster.
@@ 5.5 Ablation Studies
Here, we explore the sensitivity of the PSDE-BNN model with vertical separation of the weights to key hyperparameters. We determine the impact of the stochasticity ratio, which modulates the duration of stochastic dynamics in our framework, and the KL coefficient on model performance ($\kappa$ in (27)).
@@ 6 Conclusion and Future Work
In future work, we aim to develop an optimised numerical solver for models with horizontal cuts, enhancing their computational efficiency. We also plan to investigate models incorporating both horizontal and vertical cuts to explore potential performance improvements. Additionally, we will explore variance reduction techniques such as Sticking The Landing (Roeder et al., 2017) to improve the performance and stability of our models.
@@ Appendix A Proofs
A.1 Proof of Lemma 1
We calculate the KL-divergence as in (Xu et al., 2022). We let $w_{0}$ be a fixed initial state and $w_{1},\ldots,w_{N}$ be states at times $t_{1}=\Delta t,t_{2}=2\Delta t,\ldots,t_{N}=N\Delta t=T$ generated by the Euler discretisation:
$w_{i+1}$ $=w_{i}+f_{q}\left(t_{i},w_{i}\right)\Delta t+\sigma_{q}\left(t_{i},w_{i}\right)\left(B_{t+\Delta t}-B_{t}\right)$
$=w_{i}+f_{q}\left(t_{i},w_{i}\right)\Delta t+\sigma_{q}\left(t_{i},w_{i}\right)\Delta t^{1/2}\epsilon_{i+1}$
where $\epsilon_{i+1}\sim\mathcal{N}(0,1)$. Therefore,
$w_{i+1}\mid w_{i}\sim\mathcal{N}\left(w_{i}+f_{q}\left(t_{i},w_{i}\right)\Delta t,\sigma_{q}\left(t_{i},w_{i}\right)^{2}\Delta t\right)$
and we can write:
$\log q\left(w_{i+1}\mid w_{i}\right)=$ $-\frac{1}{2}\log\left(2\pi\sigma_{q}\left(t_{i},w_{i}\right)^{2}\Delta t\right)-\frac{1}{2}\frac{\left(w_{i+1}-\left(w_{i}+f_{q}\left(t_{i},w_{i}\right)\Delta t\right)\right)^{2}}{\sigma_{q}\left(t_{i},w_{i}\right)^{2}\Delta t}$
Similarly, for the prior process, we can write:
$\log p\left(w_{i+1}\mid w_{i}\right)=$ $-\frac{1}{2}\log\left(2\pi\sigma_{p}\left(t_{i},w_{i}\right)^{2}\Delta t\right)-\frac{1}{2}\frac{\left(w_{i+1}-\left(w_{i}+f_{p}\left(t_{i},w_{i}\right)\Delta t\right)\right)^{2}}{\sigma_{p}\left(t_{i},w_{i}\right)^{2}\Delta t}$
By substituting the expression of $w_{i+1}$ into the above equations we get:
$\log q\left(w_{i+1}\mid w_{i}\right)=$ $-\frac{1}{2}\log\left(2\pi\sigma_{q}\left(t_{i},w_{i}\right)^{2}\Delta t\right)-\frac{1}{2}\epsilon_{i+1}^{2}$
$\log p\left(w_{i+1}\mid w_{i}\right)=$ $-\frac{1}{2}\log\left(2\pi\sigma_{p}\left(t_{i},w_{i}\right)^{2}\Delta t\right)-\frac{1}{2}\Bigg{(}\frac{\left(f_{q}\left(t_{i},w_{i}\right)-f_{p}\left(t_{i},w_{i}\right)\right)^{2}}{\sigma_{p}\left(t_{i},w_{i}\right)^{2}}\Delta t$
$+\frac{2\left(f_{q}\left(t_{i},w_{i}\right)-f_{p}\left(t_{i},w_{i}\right)\right)\sigma_{q}\left(t_{i},w_{i}\right)\epsilon_{i+1}}{\sigma_{p}\left(t_{i},w_{i}\right)^{2}}\Delta t^{1/2}+\frac{\sigma_{q}(t_{i},w_{i})^{2}}{\sigma_{p}(t_{i},w_{i})^{2}}\epsilon_{i+1}^{2}\Bigg{)}$
Then, the KL-divergence for the discretised process is:
$\sum_{i=0}^{N}\mathbb{E}_{w_{i}}\left[D_{\mathrm{KL}}\left(q\left(w_{i+1}\mid w_{i}\right)\|p\left(w_{i+1}\mid w_{i}\right)\right)\right]$
$=$ $\sum_{i=0}^{N}\mathbb{E}_{w_{i}}\left[\mathbb{E}_{w_{i+1}\sim q\left(w_{i+1}\mid w_{i}\right)}\left[\log\frac{q\left(w_{i+1}\mid w_{i}\right)}{p\left(w_{i+1}\mid w_{i}\right)}\right]\right]$
$=$ $\sum_{i=0}^{N}\mathbb{E}_{w_{i}}\Bigg{[}\mathbb{E}_{\epsilon_{i+1}}\Bigg{[}A(w_{i})\Delta t+B(w_{i})\Delta t^{1/2}\epsilon_{i+1}$
$+\frac{1}{2}\left(\frac{\sigma_{q}(t_{i},w_{i})^{2}}{\sigma_{p}(w_{i})^{2}}-1\right)\epsilon_{i+1}^{2}-\frac{1}{2}\log\left(\frac{\sigma_{q}(t_{i},w_{i})^{2}}{\sigma_{p}(t_{i},w_{i})^{2}}\right)\Bigg{]}\Bigg{]}$
$=$ $\sum_{i=0}^{N}\mathbb{E}_{w_{i}}\Bigg{[}A(w_{i})\Delta t+\frac{1}{2}\left(\frac{\sigma_{q}(t_{i},w_{i})^{2}}{\sigma_{p}(t_{i},w_{i})^{2}}-1\right)-\frac{1}{2}\log\left(\frac{\sigma_{q}(t_{i},w_{i})^{2}}{\sigma_{p}(t_{i},w_{i})^{2}}\right)\Bigg{]}$
Now, taking the derivative of the function $f(x)=x^{2}-1-2\log(x)$ we can see that it has a single minimum point at $x=1$ where $f(1)=0$. Therefore, if $\sigma_{p}(t,w_{t})\neq\sigma_{q}(t,w_{t})$ with non-zero probability, then
$\mathbb{E}_{w_{t}}\left[\frac{1}{2}\left(\frac{\sigma_{q}(t,w_{t})^{2}}{\sigma_{p}(t,w_{t})^{2}}-1\right)-\frac{1}{2}\log\left(\frac{\sigma_{q}(t,w_{t})^{2}}{\sigma_{p}(t,w_{t})^{2}}\right)\right]>0$
The left-hand side is actually a continuous function of $t$ only, call it $e(t)$. Since $e(t)$ is continuous, if we take $N\rightarrow\infty$, the terms $e(t_{i})$ in the neighborhood of $t$ accumulate and make the sum blow to infinity, i.e $D_{\mathrm{KL}}\left(\mu_{q}\|\mu_{p}\right)=\infty$. ∎
@@ A.1 Proof of Lemma 1
We know that the sample path of $w_{t}$ is almost surely continuous as the Brownian motion $B_{t}$ is almost surely continuous. For any $\omega\in\Omega$ such that $w_{t}$ is continuous, we can write $w_{t}$ as a continuous function of time $t$. Therefore, given the continuous sample path $w_{t}$, we can write $f_{h}(t,h_{t};w_{t})=f_{h}(t,h_{t})$ where $f_{h}$ is continuous in $t$ and globally Lipschitz continuous in $h_{t}$. In this case, $h_{t}$ satisfies an ODE and we can follow the same argument as in Dupont et al. 2019 and show that for any $\omega$ such that $w_{t}$ is continuous, $\mathcal{L}(\psi_{1}(\omega,x))$ cannot represent $z(x)$. Actually, following the same argument we can show that $\mathcal{L}(\psi_{1}(\omega,x))$ cannot map $\{\|{x}\|\leq r_{1},x\in\mathbb{R}^{d_{x}}\}$ to the region $\{x\leq-1/2,x\in\mathbb{R}\}$ and $\{r_{2}\leq\|\mathbf{x}\|\leq r_{3},x\in\mathbb{R}^{d_{x}}\}$ to the region $\{x\geq 1/2,x\in\mathbb{R}\}$ simultaneously. If that is the case then $\psi_{1}(\{\|{x}\|\leq r_{1},x\in\mathbb{R}^{d_{x}}\})$ and $\psi_{1}(\{r_{2}\leq\|\mathbf{x}\|\leq r_{3},x\in\mathbb{R}^{d_{x}}\})$ would be linearly separable which contradicts proposition 2 in Dupont et al. 2019. Therefore, if the sample path $w_{t}$ is continuous, $\sup_{x\in A}|\mathcal{L}(\psi_{1}(\omega,x))-z(x)|>1/2$ and we have:
A.3 Proof of Theorem 4.2
First we state the noise outsourcing lemma (Kallenberg, 2010), two theorems by Zhang et al. 2020 and the theorem by Sharma et al. 2023:
@@
(Theorem 2 by Zhang et al. 2020). For any homeomorphism $h:\mathcal{X}\rightarrow\mathcal{X},\mathcal{X}\subset\mathbb{R}^{p}$, there exists a $2p$-ODE-Net $\phi_{T}:\mathbb{R}^{2p}\rightarrow\mathbb{R}^{2p}$ for $T=1$ such that $\phi_{T}\left(\left[x,0^{(p)}\right]\right)=\left[h(x),0^{(p)}\right]$ for any $x\in\mathcal{X}$.
@@
(Theorem 7 by Zhang et. al 2020). Consider a neural network $F:\mathbb{R}^{p}\rightarrow\mathbb{R}^{r}$. For $q=p+r$, there exists a linear layer-capped $q$-ODE-Net that can perform the mapping $F$.

### v2

@@ Abstract
In this paper, we present Partially Stochastic Infinitely Deep Bayesian Neural Networks, a novel family of architectures that integrates partial stochasticity into the framework of infinitely deep neural networks. Our new class of architectures is designed to improve the limitations of existing architectures around computational efficiency at training and inference time. To do this, we leverage the advantages of partial stochasticity in the infinite-depth limit which include the benefits of full stochasticity e.g. robustness, uncertainty quantification, and memory efficiency, whilst improving their limitations around computational complexity. We present a variety of architectural configurations, offering flexibility in network design including different methods for weight partition. We also provide mathematical guarantees on the expressivity of our models by establishing that our network family qualifies as Universal Conditional Distribution Approximators. Lastly, empirical evaluations across multiple tasks show that our proposed architectures achieve better downstream task performance and uncertainty quantification than their counterparts while being significantly more efficient.
@@ 1 Introduction
To overcome these challenges, previous research has explored partially stochastic networks as a potential solution in finite-depth settings (Daxberger et al. 2021b; Kristiadi et al. 2020; Li et al. 2024). These networks restrict stochasticity to a fraction of the model’s parameters, i.e. $\Theta=\Theta_{S}\cup\Theta_{D}$, where one learns both a distribution over $\Theta_{S}$ and point estimates for $\Theta_{D}$. These methods aim to achieve greater computational efficiency while retaining the benefits of Bayesian approaches. Initially, the main idea of these architectures was to approximate fully stochastic networks with a trade-off between computational efficiency and accuracy. However, Sharma et al. (2023) provide both theoretical and empirical evidence to demonstrate that partially stochastic networks can achieve comparable, if not superior, performance to their fully stochastic counterparts. While Sharma et al. (2023) leaves the selection of non-stochastic components in BNNs largely unexplored, works around Variational Bayesian Last Layers like Harrison et al. (2024), address this issue. As there is no trivial way to split the weights into subsets of deterministic and stochastic weights in the infinite-depth limit, previous research has only explored finite-depth regimes. However, studies such as Farquhar et al. (2020) hint towards the viability of extending this approach to an infinite-depth configuration.
@@ 2 Background and Preliminaries
In Bayesian Neural Networks (Lampinen & Vehtari, 2001; Titterington, 2004), parameters are treated as random variables, with the aim to infer their posterior distribution $p(w|\mathcal{D})$ given a dataset $\mathcal{D}$ and prior $p(w)$. Approximate inference involves finding an approximate posterior, often through a parametrisation $q_{\theta}(w)$. Specifically, variational inference methods intend to maximise the Evidence Lower Bound (ELBO), or equivalently, minimise the Kullback-Leibler (KL) divergence between $q_{\theta}(w)$ and $p(w|\mathcal{D})$. The gradient of this objective can be computed using Monte-Carlo samples via Stochastic Variational Inference (Hoffman et al., 2013). In addition, previous studies suggest that full stochasticity is unnecessary for effective posterior approximation in finite-depth BNNs (Sharma et al., 2023). Extending these findings, we show that partial stochasticity preserves the expressivity of BNNs in the infinite-depth regime.
@@
where $\mu:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ (the drift), $\sigma:[0,T]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times m}$ (the volatility or diffusion matrix) are Lipschitz functions in the state, and satisfy a linear growth condition. Such conditions ensure the existence and uniqueness of a strong solution, given a square-integrable initial condition $X_{0}$. SDEs are a natural way of embedding stochasticity in a continuous-time framework, hence they are fit for modeling infinitely deep networks in a Bayesian setting. In infinite-depth Bayesian neural networks, Xu et al. (2022) propose to model the prior on weights $(w_{t})_{t\in[0,1]}$ as a SDE
@@
and parametrize the approximate posterior using a Neural SDE, where the drift $f_{q}$ is a neural network with parameters $\theta$
@@
Variational inference on SDEs was introduced by Archambeau et al. (2007), and previous work have relied on other techniques to learn the coefficients of SDEs, such as Karhunen-Loève expansion of Brownian motion (Ghosh et al., 2022). The parametrization of the weights given by (2) require using KL divergence on path spaces. Minimizing the KL divergence between path measures induced by diffusion processes is a well-studied mathematical problem (Cattiaux & Léonard, 1994) which has been recently leveraged in the context of learning the Schrödinger Bridge Problem (Vargas et al., 2023) and regularizing generative models (Richter et al., 2024; Vargas et al., 2024). We use the following infinite-dimensional version of the KL divergence (Tzen & Raginsky, 2019; Li et al., 2020)
@@
with $\mu_{p},\mu_{q}$ being the prior and posterior measures on the path space $\mathcal{P}\subset\mathbb{R}^{[0,1]}$ induced by (1) and (2) respectively. We show that when only a fraction of the weights are stochastic, such approximate posteriors yield UCDAs.
@@ 3 Method
where $t$ is the layer index, $h_{t}\in\mathbb{R}^{d_{h}}$ represents a vector of hidden unit activations at layer $t$, the input is $h_{0}=x$, $w_{t}$ are the parameters for layer t, and $\delta=1$. Taking $\delta=1/T$ and letting $T\rightarrow\infty$ yields a Neural ODE $\mathrm{d}h_{t}=f_{h}(t,h_{t};w_{t})\mathrm{d}t$, where $f_{h}$ is a neural network with parameters $w_{t}$. These parameters can also be modelled by a neural network (more precisely, a hypernetwork), which in the infinite-depth limit becomes itself a neural differential equation $\mathrm{d}w_{t}=f_{q}(t,w_{t};\theta)\mathrm{d}t$ (or, in the Bayesian setting, a Neural SDE). Therefore, $f_{q}$ is a neural network parametrised by $\theta$ which models the continuous evolution of the parameters of $f_{h}$. Finally, in this framework, the only trainable parameters are the vector $\theta$ and $w_{0}$. While we use SDE formalism similar to generative diffusion models, our model is discriminative. It ultimately learns the conditional distribution $p(y\mid x)$ via continuous latent variables.
@@ 3.1 Vertical separation of the weights
Notice that the diffusion of the approximate posterior is fixed to match the prior diffusion. This is to ensure that the KL divergence between the prior and posterior measures on paths $\mu_{p}$ and $\mu_{q}$ respectively, is finite. The proof of the following lemma can be found in (Archambeau et al., 2007).
@@
In conclusion. the dynamics of the whole system can then be written as a single SDE in $\mathbb{R}^{d_{h}+d_{w}}$:
@@ 4.2 Partially Stochastic Infinitely Deep Bayesian Neural Networks are UCDAs
ResNet32 99.46 $\pm$ 0.00 2.88 ±$\pm 0.94$ 87.35 ±$\pm 0.00$ 8.47 $\pm$ 0.39
ODENet 98.90 ±$\pm 0.04$ 1.11 $\pm$ 0.10 88.30 $\pm$ 0.29 8.71 ±$\pm 0.21$
MFVI ResNet 32 99.40 $\pm$ 0.00 2.76 $\pm$ 1.28 86.97 $\pm$ 0.00 3.04 $\pm$ 0.94
MFVI 86.48 1.95
@@
Deep Ensemble 89.22 2.79
HMC (“gold standard”) 98.31 1.79 90.70 5.94
MFVI ODENet 98.81 ±$\pm 0.00$ 2.63 ±$\pm 0.31$ 81.59 ±$\pm 0.01$ 3.62 ±$\pm 0.40$
MFVI HyperODENet 98.77 ±$\pm 0.01$ 2.82 ±$\pm 1.34$ 80.62 ±$\pm 0.00$ 4.29 ±$\pm 1.10$
@@ 5.1 Experimental Setup
We conduct our experiments under three distinct configurations. The first configuration, termed ‘SDEfirst’, is defined by setting the initial time cut, $t_{1}$, to 0. In this setup, the weights process is initially modelled with a SDE and then transitions to an ODE regime. Within this framework, we explore two sub-settings: one where the initial values of the ODE are the weights obtained from the SDE, and another, labelled ‘fix_w2’, where the initial values of the ODE are trainable parameters. The second configuration, ‘ODEFirst’, sets the second time cut, $t_{2}$, to 1, i.e. the endpoint of the considered time frame. Here, the weight evolution initially follows an ODE, with stochasticity introduced in the subsequent phase. The last configuration, the Horizontal Cut method (‘Hor. Cut’), splits the weight vector into two subsets, $S$ and $D$ here $t_{1}=0$ and $t_{2}=1$. These segments evolve according to an SDE, discretised using the Euler-Maruyama scheme, and an ODE computed using the Midpoint scheme, respectively. We emphasize two essential criteria on choosing $t_{1}$ and $t_{2}$. First is the data’s uncertainty and complexity. For data characterized by high uncertainty, or limited prior knowledge, extending the stochastic phase ($t_{2}$ in SDEFirst) may be beneficial. Conversely, when more prior information is available, prolonging the deterministic phase ($t_{1}$ in ODEFirst) could be more beneficial for the model. Second is the computational efficiency. Models employing a longer stochastic phase could demand greater computational resources, a factor that needs attention in scenarios with limited computational budgets. One can also refer to table 4 to see the impact of $t_{1}$ and $t_{2}$ on the performance and computing time of our models.
For hyperparameter optimisation, we employ a comprehensive grid search, which is detailed in Appendix E. Across all experiments, the diffusion function $g_{p}$ is set to a constant value. Lastly, experiments were conducted on a single Nvidia RTX 3090 GPU.
@@ 5.3 Uncertainty Quantification
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Model Corruption Lvl 2 Corruption Lvl 3 Corruption Lvl 4 Corruption Lvl 5
ODEFirst $0.322\pm 0.409$ $0.381\pm 0.431$ $0.431\pm 0.450$ $0.503\pm 0.465$
SDEFirst $\mathbf{0.443\pm 0.468}$ $\mathbf{0.498\pm 0.480}$ $\mathbf{0.543\pm 0.489}$ $\mathbf{0.628\pm 0.512}$
Hor. Cut $0.394\pm 0.436$ $0.430\pm 0.447$ $0.488\pm 0.452$ $0.541\pm 0.458$
SDEBNN $0.256\pm 0.353$ $0.293\pm 0.370$ $0.333\pm 0.388$ $0.375\pm 0.405$
Table 2: Quantitative evaluation of model performance under varying corruption levels in the CIFAR10-C dataset. All models exhibit an increase in predictive entropy as the level of corruption intensifies, demonstrating their sensitivity to image degradation. Notably, models such as SDEBNN display a more pronounced increase in entropy across corruption levels, indicating a robust capability to recognize and respond to deteriorating input quality.
In the discrete neural networks domain, partial stochasticity is initially conceptualised as a simplified approximation of fully stochastic neural networks aimed at providing just enough uncertainty quantification. However, as we have theoretically established in earlier sections, partial stochasticity provides an expressive framework to approximate any posterior distribution. In this section, we further prove this empirically by looking at the expected calibration error (ECE) of the models at test time, and by performing predictive entropy analysis of out-of-distribution (OOD) and in-distribution (ID) samples.
Table 1 shows that across the uncertainty-aware models, our PSDE-BNN (SDEFirst) surpasses all baseline models in ECE performance. While ODEFirst, SDEFirst with ‘fix_w2’, and PSDE-BNN (Hor. Cut) consistently demonstrate competitive error calibration and accuracy, SDE-BNN faces challenges in maintaining low ECE when striving for higher accuracy. This is because, during training, a trade-off emerges: ECE is sacrificed for improved accuracy, likely due to overfitting that narrows uncertainty estimates over time. The phenomenon is accentuated for the SDE-BNN which, at relatively low accuracies for CIFAR-10 (in the range of 70% - 80%) can achieve good ECE performance which quickly - and at a much faster rate than PSDE-BNNs - decays over epochs.
@@
Additionally, we have used CIFAR10-C (Hendrycks & Dietterich, 2019), a dataset which introduces various corruption levels to systematically assess our models’ uncertainty estimation. Our investigations included adding 17 different types of corruptions (e.g. gaussian noise, blur, snowification, compression, pixelation, etc) at 5 different levels of intensity to challenge the models’ robustness and generalisation. Results are shown in Table 2.
@@ 5.4 Efficiency and Learning Rate
Table 3: Models’ average times in seconds.
Model Epoch Time Inference Time
@@
ODEFirst (10%) 289.4 31.7
SDEFirst (10%) $295.1$ 31.7
Hor. Cut (50%) $316.7$ 35.9
Next, we explore the learning rate of our models, i.e. how many epochs the network takes to get to acceptable downstream task performance. PSDE-BNN takes considerably less time to achieve the reported classification accuracies. While our models trained for 100 epochs, SDE-BNN (and as reported by Xu et al. (2022)) required 300 epochs. The evolution after the 80% test accuracy for SDE-BNN was especially long. This indicates that PSDE-BNNs were both faster at each epoch, and only needed a third of the epochs to achieve similar (although better) downstream performance. Therefore, the overall training time was 74.1% faster.
@@ 5.5 Ablation Studies
Here, we explore the sensitivity of the PSDE-BNN model with a vertical separation of the weights to key hyperparameters. We determine the impact of the stochasticity ratio, which modulates the duration of stochastic dynamics in our framework, and the KL coefficient ($\kappa$ in (12)) on model performance.
@@ 6 Conclusion and Future Work
In future work, we aim to investigate models incorporating both horizontal and vertical cuts to explore potential performance improvements. Additionally, we will explore variance reduction techniques such as Sticking The Landing (Roeder et al., 2017) to improve the performance and stability of our models. Furthermore, an interesting direction to evaluate the generalization power of our approaches is to test them on a variety of datasets, including but not limited to Omniglot for few-shot learning and ImageNet for performance in high-resolution datasets. As scalability is a common issue with Bayesian models, implementing PSDE-BNN on learned features might help further improve the efficiency of the model. At last, it is relevant to investigate the relationship between the test set performance and the quality of the variational approximation more systematically, including exploring different priors and model complexities to better understand how the KL divergence impacts generalization across various settings.
@@ Appendix A Proofs
A.1 Proof of Theorem 4.1
We know that the sample path of $w_{t}$ is almost surely continuous as the Brownian motion $B_{t}$ is almost surely continuous. For any $\omega\in\Omega$ such that $w_{t}$ is continuous, we can write $w_{t}$ as a continuous function of time $t$. Therefore, given the continuous sample path $w_{t}$, we can write $f_{h}(t,h_{t};w_{t})=f_{h}(t,h_{t})$ where $f_{h}$ is continuous in $t$ and globally Lipschitz continuous in $h_{t}$. In this case, $h_{t}$ satisfies an ODE and we can follow the same argument as in Dupont et al. (2019) and show that for any $\omega$ such that $w_{t}$ is continuous, $\mathcal{L}(\psi_{1}(\omega,x))$ cannot represent $z(x)$. Actually, following the same argument we can show that $\mathcal{L}(\psi_{1}(\omega,x))$ cannot map $\{\|{x}\|\leq r_{1},x\in\mathbb{R}^{d_{x}}\}$ to the region $\{x\leq-1/2,x\in\mathbb{R}\}$ and $\{r_{2}\leq\|\mathbf{x}\|\leq r_{3},x\in\mathbb{R}^{d_{x}}\}$ to the region $\{x\geq 1/2,x\in\mathbb{R}\}$ simultaneously. If that is the case then $\psi_{1}(\{\|{x}\|\leq r_{1},x\in\mathbb{R}^{d_{x}}\})$ and $\psi_{1}(\{r_{2}\leq\|\mathbf{x}\|\leq r_{3},x\in\mathbb{R}^{d_{x}}\})$ would be linearly separable which contradicts proposition 2 in Dupont et al. (2019). Therefore, if the sample path $w_{t}$ is continuous, $\sup_{x\in A}|\mathcal{L}(\psi_{1}(\omega,x))-z(x)|>1/2$ and we have:
@@
First we state the noise outsourcing lemma (Kallenberg, 2010), two theorems by Zhang et al. (2020) and the theorem by Sharma et al. (2023):
@@
(Theorem 2 by Zhang et al. (2020)). For any homeomorphism $h:\mathcal{X}\rightarrow\mathcal{X},\mathcal{X}\subset\mathbb{R}^{p}$, there exists a $2p$-ODE-Net $\phi_{T}:\mathbb{R}^{2p}\rightarrow\mathbb{R}^{2p}$ for $T=1$ such that $\phi_{T}\left(\left[x,0^{(p)}\right]\right)=\left[h(x),0^{(p)}\right]$ for any $x\in\mathcal{X}$.
@@
(Theorem 7 by Zhang et. al (2020)). Consider a neural network $F:\mathbb{R}^{p}\rightarrow\mathbb{R}^{r}$. For $q=p+r$, there exists a linear layer-capped $q$-ODE-Net that can perform the mapping $F$.